Find a polynomial for the perimeter and the area.
one side is = b+2
the other side is= b
The perimeter is______(simplify your answer. do not factor)
The area is_______(simplify your answer. do not factor)

Answer by checkley77(12569)

one side is = b+2
the other side is= b
Perimeter=2L+2W
PERIMETER=2(B+2)+2B
PERIMETER=2B+4+2B
PERIMETER=4B+4 OR 4(B+1) ANS.
AREA=LW
AREA=B(B+2) ANS.